# Cooperative Sheaf Neural Networks

### v1

@@ 4 Cooperative Sheaf Neural Networks
If the target map of $i$ is zero, then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{i}=0$. If ${\mathbf{S}}_{k}=0$ for some neighbor $k$ of $i$, then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{i}$ does not depend on $\mathbf{x}_{k}$.
@@
Moreover, as a consequence of Equation 7, our model has the ability to reach longer distances:
Proposition 4.2.
@@
Let $t=1$, and fix a node $i$. Then we are essentially just using the composition described in Equation 7 (up to normalization and learnable weights). In the equation we have a sum running over all neighbors $j$ of $i$ and another sum running over all neighbors $u$ of each $j$. So $u$ can be a $2$-hop neighbor of $i$ and we have that $i$ was updated with information from up to $2$-hop neighbors. Similarly, the node $u$ is updated by up to $2$-hop neighbors. Therefore, in the second layer $t=2$, $i$ was updated with information from up to $4$-hop neighbors.
@@
Proposition 4.2 shows that CSNN can reach longer distances: in most GNNs if $t$ is the distance between two nodes $i$ and $j$, then they can only communicate after $t$ layers. CSNN enables communication between these nodes after $\lceil t/2\rceil$ layers. Moreover, the following result shows that CSNNs with $t$ layers are capable of making $i$ and $j$ communicate while ignoring all the other nodes on a path from $i$ to $j$, as proved in the next result.
Proposition 4.3.
Let $i$ and $j$ be nodes at a distance $t$. In CSNN, $i$ can learn to ignore all the $t-1$ nodes in the shortest path from $i$ to $j$ while receiving the information from $j$ in the $t$-layer. Moreover, if we choose a path with $n>t-1$ nodes between $i$ and $j$, then $i$ receives the information from $j$ in the $(n+1)$-layer.
@@
Choose a path from $i$ to $j$. So there are $t-1$ vertices between $i$ an $j$, say $v_{1},...,v_{t-1}$. In the first layer, let ${\mathbf{S}}_{j}$ and ${\mathbf{T}}_{v_{t-1}}$ be different of zero and all other source and target maps equal zero.
@@
The above shows that by learning the right sheaves in each layer, if a node $i$ wants to receive information only from node $j$ at distance $t$, then it can ignore all other nodes until it reaches the $t$-layer, where it will get the information from node $j$ that was updated only by itself, layer by layer. Intuitively, a node $i$ will be updated by a neighbor $k$ only if it is able do listen and $k$ is able to broadcast. This is the setup we use to go from one node to the next one without updating anything else until it reaches the final node. Although we need all the $t$-layers, this is an advantage compared to standard message-passing since we have the possibility to ignore other nodes.
Example 4.4.
Consider a graph formed by vertices $\{1,2,3,4,5\}$ and edges $\{12,23,34,45,54,43,32,21\}$. We follow the proof of 4.3 to check we can propagate a message from $5$ to $1$, while $1$ ignores all other nodes. In the first layer we must have that all source and target maps are zero except by ${\mathbf{T}}_{4}$ and ${\mathbf{S}}_{5}$. A simple verification gives that $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{k}=0$, for all $k\neq 4$. So $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{k}=0$, for all $k\neq 4$. If $k=4$, then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{4}=-2{\mathbf{T}}_{4}^{\top}{\mathbf{S}}_{5}\textbf{x}_{5}^{(0)},$ with $\textbf{x}_{k}^{(t)}$ denoting the feature vector of $k$ at layer $t$, thus $\textbf{x}_{5}^{(0)}$ will be the only feature vector to influence $\textbf{x}_{4}^{(1)}$.
In the second layer we must have that all source and target maps are zero except by ${\mathbf{T}}_{3}$ and ${\mathbf{S}}_{4}$. Then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{k}=0$, for all $k\neq 3$ and $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{3}=-2{\mathbf{T}}_{3}^{\top}{\mathbf{S}}_{4}\textbf{x}_{4}^{(1)}$. Thus $\textbf{x}_{4}^{(1)}$ will be the only feature vector to influence $\textbf{x}_{3}^{(2)}$.
In the third layer we must have that all source and target maps are zero except by ${\mathbf{T}}_{2}$ and ${\mathbf{S}}_{3}$. Then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{k}=0$, for all $k\neq 2$ and $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{2}=-2{\mathbf{T}}_{2}^{\top}{\mathbf{S}}_{3}\textbf{x}_{3}^{(3)}$. Thus $\textbf{x}_{3}^{(2)}$ will be the only feature vector to influence $\textbf{x}_{2}^{(3)}$.
In the fourth and last layer we must have that all source and target maps are zero except by ${\mathbf{T}}_{1}$ and ${\mathbf{S}}_{2}$. Then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{k}=0$, for all $k\neq 1$ and $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{1}=-2{\mathbf{T}}_{1}^{\top}{\mathbf{S}}_{2}\textbf{x}_{2}^{(3)}$. Thus $\textbf{x}_{1}^{(4)}$ is influence only by $\textbf{x}_{2}^{(3)}$, which was influenced only by $\textbf{x}_{3}^{(2)}$, which in turn was influenced only by $\textbf{x}_{4}^{(1)}$, which receive information only from $\textbf{x}_{5}^{(0)}$. Consequently, $\textbf{x}_{1}^{(4)}$ is affected by $\textbf{x}_{5}^{(0)}$ while ignoring the feature vectors of all other nodes in all other layers.
Figure 3: Illustration of Example 4.4. At layer $\ell$, we consider that all maps but ${\mathbf{T}}_{5-\ell}$ and ${\mathbf{S}}_{5-(\ell-1)}$ are 00, enabling the flow of information from right to left following the bottom edges.
Observe that the derivative of $\textbf{x}_{1}^{(4)}$ in relation to $\textbf{x}_{5}^{(0)}$ can be as high as the values of the non-zero ${\mathbf{T}}_{i}$ and ${\mathbf{S}}_{i}$ permit. This shows our model can mitigate over-squashing, which refers to the failure of an information propagating to distance nodes. Di Giovanni et al. (2023) and Topping et al. (2022), studied over-squashing in message passing neural networks through a bound on the Jacobian
$\left|\frac{\partial\textbf{x}_{i}^{(t)}}{\partial\textbf{x}_{j}^{(0)}}\right|\leq c^{t}\hat{A}^{t}_{ij},$ (8)
where $t$ is the layer, $c$ is a constant that depends on the architecture of the model, and $\hat{A}$ the normalized adjacency matrix. Moreover, over-squashing occurs when we have a small derivative $\nicefrac{{\partial\textbf{x}_{i}^{(t)}}}{{\partial\textbf{x}_{j}^{(0)}}}$, since it means that after $t$ layers, the feature at $i$ is mostly insensitive to the information initially contained at $j$, i.e., the information was not propagated properly. Proposition 4.3 says that the feature at $i$ can be sensitive to the information initially contained at $j$, independently of the distance, if the right sheaves are learned and sufficient layers are used.
5 Experimental Results
This section provides real-world experiments to validate the proposed model. We use the recently proposed benchmarks from Platonov et al. (2023) to test our model capacity on dealing with node classification tasks focusing on heterophilic settings, which are notoriously challenging for GNNs. Additionally, we check the capacity of CSNN to deal with over-squashing on a synthetic experiment.
5.1 Over-squashing
In order to verify our theoretical results on the capacity of CSNN to alleviate over-squashing, we reproduce the NeighborsMatch problem proposed by Alon and Yahav (2020), using the same framework. The datasets consist of binary trees of depth $r$, with the root node as the target, the leaves containing its possible labels, and the leaf with the same number of neighbors as the target node containing its true label. We provide the parameters used for this task in Appendix B.
Figure 4 shows that GCN and GIN fail to fit the datasets starting from $r=4$ and GAT and GGNN fail to fit the datasets starting from $r=5$. These models suffered from over-squashing and were not able to distinguish between different training examples, while the CSNN model reached perfect accuracy for all tested $r$.

### v2

@@ 4.1 Analysis
If the target map ${\mathbf{T}}_{i}$ is zero, then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{i}=0$. If the source map ${\mathbf{S}}_{k}=0$ for some neighbor $k$ of $i$, then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{i}$ does not depend on $\mathbf{x}_{k}$.
Moreover, our model has the ability to reach longer distances. In most GNNs, if $t$ is the distance between two nodes $i$ and $j$, then they can only communicate after $t$ layers. CSNN enables communication between these nodes after $\lceil t/2\rceil$ layers.
Proposition 4.2.
We also show in Proposition˜4.3 that CSNNs with $t$ layers are capable of making $i$ and $j$ communicate while ignoring all the other nodes on a path from $i$ to $j$ such that $|j-i|\leq t$. This feature is an asset to handle over-squashing in long range tasks, allowing CSNN to selectively tend to information from distant nodes. Example˜4.4 illustrates this result in a four-node graph.
Proposition 4.3.
Let $i$ and $j$ be nodes at a distance $t$. In CSNN, $i$ can learn to ignore all the t1t-1 nodes in the shortest path from $i$ to $j$ while receiving the information from $j$ in the $t$-layer. Moreover, if we choose a path with n>t1n>t-1 nodes between $i$ and $j$, then $i$ receives the information from $j$ in the $(n+1)$-layer.
Example 4.4.
Consider a directed graph with vertex set $V=\{1,2,3,4\}$ and edge set $E=\{(1,2),$ (2,3),(3,4),(4,3),(3,2),(2,1)}(2,3),(3,4),(4,3),(3,2),(2,1)\}. We follow the proof of Proposition˜4.3 to show we can propagate a message from node $4$ to node $1$, while the latter ignores all remaining nodes. We denote the target/source map of node $i$ at layer $t$ by ${\mathbf{T}}_{i,t}$ and ${\mathbf{S}}_{i,t}$ respectively. To achieve our desired result:
  1. a.
    In the first layer, must have zero source and target maps except for ${\mathbf{T}}_{3,1}$ and ${\mathbf{S}}_{4,1}$. A simple verification gives that $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{k}=0$, for all $k\neq 3$. So $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{k}=0$, for all $k\neq 3$. If $k=3$, then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{3}=-2{\mathbf{T}}_{4,1}^{\top}{\mathbf{S}}_{4,1}\mathbf{x}_{4}^{(0)},$ with $\mathbf{x}_{k}^{(t)}$ denoting the feature vector of $k$ at layer $t$, thus $\mathbf{x}_{4}^{(0)}$ will be the only feature vector to influence $\mathbf{x}_{3}^{(1)}$;
  2. b.
    In the second layer, we must have that all source and target maps are zero except for ${\mathbf{T}}_{2,2}$ and ${\mathbf{S}}_{3,2}$. Then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{k}=0$, for all $k\neq 2$ and $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{2}=-2{\mathbf{T}}_{2,2}^{\top}{\mathbf{S}}_{3,2}\mathbf{x}_{3}^{(1)}$. Thus $\mathbf{x}_{3}^{(1)}$ will be the only feature vector to influence $\mathbf{x}_{2}^{(2)}$;
  3. c.
    In the third layer, all source and target maps must be zero except for ${\mathbf{T}}_{1,3}$ and ${\mathbf{S}}_{2,3}$. Then $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{k}=0$, for all $k\neq 1$ and $((L^{\text{in}}_{\mathcal{F}})^{\top}L^{\text{out}}_{\mathcal{F}}(\mathbf{X}))_{1}=-2{\mathbf{T}}_{1,3}^{\top}{\mathbf{S}}_{2,3}\mathbf{x}_{2}^{(3)}$. Thus $\mathbf{x}_{1}^{(3)}$ is influenced only by $\mathbf{x}_{2}^{(2)}$, which was influenced only by $\mathbf{x}_{3}^{(1)}$, which is influenced only by $\mathbf{x}_{4}^{(0)}$.
Consequently, $\mathbf{x}_{1}^{(3)}$ is affected by $\mathbf{x}_{4}^{(0)}$ while ignoring the features of all other nodes in all other layers.
Figure 3: Illustration of Example˜$4.$ At layer $t$, we consider that all maps but ${\mathbf{T}}_{4-t,t}$ and ${\mathbf{S}}_{4-(t-1),t}$ are 0, enabling the flow of information from right to left following the bottom edges.
Observe that the derivative of $\mathbf{x}_{1}^{(3)}$ in relation to $\mathbf{x}_{4}^{(0)}$ can be as high as the values of the non-zero ${\mathbf{T}}_{i}$ and ${\mathbf{S}}_{i}$ permit. This shows our model can mitigate over-squashing, which refers to the failure of an information propagating to distance nodes. Di Giovanni et al. (2023) and Topping et al. (2022), studied over-squashing in message passing neural networks through a bound on the Jacobian $\left|\nicefrac{{\partial\mathbf{x}_{i}^{(t)}}}{{\partial\mathbf{x}_{j}^{(0)}}}\right|\leq c^{t}\hat{A}^{t}_{ij},$ where $t$ is the layer, $c$ is a constant that depends on the architecture of the model, and $\hat{A}$ the normalized adjacency matrix. Moreover, over-squashing occurs when we have a small derivative $\nicefrac{{\partial\mathbf{x}_{i}^{(t)}}}{{\partial\mathbf{x}_{j}^{(0)}}}$, since it means that after $t$ layers, the feature at $i$ is mostly insensitive to the information initially contained at $j$, i.e., the information was not propagated properly. Proposition˜4.3 states that the feature at node $i$ can be sensitive to the information initially contained at node $j$, independently of the distance, given enough layers and an appropriate configuration of restriction maps.
5 Related works
Cooperative GNNs. Finkelshtein et al. (2024) were the first to propose flexibilizing message passing by allowing nodes to choose how they cooperate with each other. Each layer of their model, CO-GNN, employs an additional GNN that chooses the cooperation the action each node takes. While CO-GNNs can be employed with arbitrary base and action networks, their main caveat is that training can become increasingly difficult as these networks become more complex — the grid of hyper-parameters grows considerably and the stochastic nature of the action network may affect model selection. Different from CO-GNN, our CSNN does not rely on discrete actions and can smoothly modulate between cooperative behavior patterns.
Sheaf Neural Networks. Besides works on SNNs for graph data with real-valued node features, recent works have also expanded the literature to deal with heterogenous edge types  (Braithwaite et al., 2024), hypergraphs (Duta et al., 2023), and node features living on Riemann manifolds  (Battiloro et al., 2024). Additionally, while recent works on SNNs learn restriction maps in an end-to-end fashion, there are also prior works in which they are manually constructed  (Hansen and Ghrist, 2019) or computed as a pre-processing step (Barbero et al., 2022).
6 Experiments
We provide both synthetic and real-world experiments to evaluate the performance of CSNN, including node- and graph-level prediction tasks. Section˜6.1 assess CSNN’s capacity to circumvent over-squashing using the NeighborsMatch benchmark proposed by Alon and Yahav (2021). Section˜6.2 presents experiments on eleven node classification tasks, showcasing the effectiveness of CSNN for heterophilic graphs. Finally, Section˜6.3 consider the Peptides datasets from the Long Range Graph Benchmark (Dwivedi et al., 2022) to substantiate the capability of our model to mitigate under-reaching and over-squashing on real-world graph classification.
6.1 Over-squashing
In order to verify our theoretical results on the capacity of CSNN to alleviate over-squashing, we reproduce the NeighborsMatch problem proposed by Alon and Yahav (2021), using the same framework. The datasets consist of binary trees of depth $r$, with the root node as the target, the leaves containing its possible labels, and the leaf with the same number of neighbors as the target node containing its true label. We provide the parameters used for this task in Appendix˜B.
Figure 4: Accuracy for increasing tree depths in the NeighborsMatch task. CSNN consistently achieves 100% accuracy for all values of $r$.
This experiment shows that CSNN is more efficient in ignoring irrelevant nodes and can avoid loosing relevant information. Moreover, Proposition˜4.3 provides theoretical support for this result, as it states that there are choices of parameters for which the model can listen only to the nodes along a path between distant nodes $i$ and $j$, enabling selective communication to diminish noise impact.
Figure˜4 shows GCN (Kipf and Welling, 2017) and GIN (Xu et al., 2019) fail to fit the datasets starting from $r=4$ and GAT (Veličković et al., 2018) and GGNN (Li et al., 2016) fail to fit the datasets starting from $r=5$. These models suffer from over-squashing and are not able to distinguish between different training examples, while the CSNN model reaches perfect accuracy for all tested $r$.
Alon and Yahav (2021) argue that the difference in performance for the GNNs are related to how node features are updated: on one hand, GCN and GIN aggregate all neighbor information before combining it with the representation of the target node, forcing them to compress all incoming information into a single vector. On the other hand, GAT uses attention to selectively weigh messages based on the representation of the target node, allowing it (to some degree) to filter out irrelevant edges and only compress information from a subset of the neighbors. So models like GAT (and GGNN) that compress less information per step can handle higher $r$ better than GCN and GIN.
6.2 Node Classification
Datasets. We evaluate our model on the five recently proposed heterophilic graphs from Platonov et al. (2023), and also on six classical ones for which benchmarking results can be found in Pei et al. (2020); Rozemberczki et al. (2021); Tang et al. (2009). As pointed out in Platonov et al. (2023), the datasets Squirrel and Chameleon have many duplicate nodes, which may lead to data leakage. Following their guidelines, we use their cleaned version of these datasets to ensure a meaningful evaluation. For binary classification datasets, we report AUROC, while for multiclass datasets we report accuracy.
Table 1: Performance comparison on datasets from Platonov et al. (2023). AUROC is reported for minesweeper, tolokers and questions, accuracy is reported for the remaining datasets. CSNN is the best-performing method in 6 out of 7 datasets.
Model roman-empire amazon-ratings minesweeper tolokers questions squirrel chameleon
GCN 73.69 ± 0.74 48.70 ± 0.63 89.75 ± 0.52 83.64 ± 0.67 76.09 ± 1.27 39.47 ± 1.47 40.89 ± 4.12
SAGE 85.74 ± 0.67 53.63 ± 0.39 93.51 ± 0.57 82.43 ± 0.44 76.44 ± 0.62 36.09 ± 1.99 37.77 ± 4.14
GAT 80.87 ± 0.30 49.09 ± 0.63 92.01 ± 0.68 83.70 ± 0.47 77.43 ± 1.20 35.62 ± 2.06 39.21 ± 3.08
GAT-sep 88.75 ± 0.41 52.70 ± 0.62 93.91 ± 0.35 83.78 ± 0.43 76.79 ± 0.71 35.46 ± 3.10 39.26 ± 2.50
GT 86.51 ± 0.73 51.17 ± 0.66 91.85 ± 0.76 83.23 ± 0.64 77.95 ± 0.68 36.30 ± 1.98 38.87 ± 3.66
GT-sep 87.32 ± 0.39 52.18 ± 0.80 92.29 ± 0.47 82.52 ± 0.92 78.05 ± 0.93 36.66 ± 1.63 40.31 ± 3.01
CO-GNN 89.44 ± 0.50 54.20 ± 0.34 97.35 ± 0.63 84.84 ± 0.96 75.97 ± 0.89 39.39 ± 2.76 41.14 ± 5.40
O(d)-NSD 80.41 ± 0.72 42.76 ± 0.54 92.15 ± 0.84 78.83 ± 0.76 69.69 ± 1.46 35.79 ± 3.34 37.93 ± 2.24
BuNN 91.75 ± 0.39 53.74 ± 0.51 98.99 ± 0.16 84.78 ± 0.80 78.75 ± 1.09 - -
CSNN 92.63 ± 0.50 52.07 ± 1.00 99.07 ± 0.25 85.45 ± 0.53 79.31 ± 1.22 41.18 ± 2.23 43.09 ± 3.17
Table 2: Accuracy for node classification datasets on the fixed splits of Pei et al. (2020). CSNN achieves the best results in 3 out of 4 datasets.
Model Texas Wisconsin Film Cornell
GGCN 84.86 ± 4.55 86.86 ± 3.29 37.54 ± 1.56 85.68 ± 6.63
H2GCN 84.86 ± 7.23 87.65 ± 4.98 35.70 ± 1.00 82.70 ± 5.28
GPRGNN 78.38 ± 4.36 82.94 ± 4.21 34.63 ± 1.22 80.27 ± 8.11
FAGCN 82.43 ± 6.89 82.94 ± 7.95 34.87 ± 1.25 79.19 ± 9.79
MixHop 77.84 ± 7.73 75.88 ± 4.90 32.22 ± 2.34 73.51 ± 6.34
GCNII 77.57 ± 3.83 80.39 ± 3.40 37.44 ± 1.30 77.86 ± 3.79
Geom-GCN 66.76 ± 2.72 64.51 ± 3.66 31.59 ± 1.15 60.54 ± 3.67
PairNorm 60.27 ± 4.34 48.43 ± 6.14 27.40 ± 1.24 58.92 ± 3.15
GraphSAGE 82.43 ± 6.14 81.18 ± 5.56 34.23 ± 0.99 75.95 ± 5.01
GCN 55.14 ± 5.16 51.76 ± 3.06 27.32 ± 1.10 60.54 ± 5.30
GAT 52.16 ± 6.63 49.41 ± 4.09 27.44 ± 0.89 61.89 ± 5.05
MLP 80.81 ± 4.75 85.29 ± 3.31 36.53 ± 0.70 81.89 ± 6.40
Diag-NSD 85.67 ± 6.95 88.63 ± 2.75 37.79 ± 1.01 86.49 ± 7.35
O(d)-NSD 85.95 ± 5.51 89.41 ± 4.74 37.81 ± 1.15 84.86 ± 4.71
Gen-NSD 82.97 ± 5.13 89.21 ± 3.84 37.80 ± 1.22 85.68 ± 6.51
CSNN 87.30 ± 5.93 90.00 ± 2.83 38.03 ± 1.12 81.62 ± 4.32
Experimental setting. As baselines for benchmarks in Platonov et al. (2023), we use GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018) and GT (Shi et al., 2021), together with the variations GAT-sep and GT-sep, which concatenate the representation of a node to the mean of its neighbors instead of summing them (Zhu et al., 2020). These are all classical baselines used in Platonov et al. (2023) to compare against GNN architectures specifically developed for heterophilic settings, and that achieve the best performance in most cases. We also compare CSNN against recent models such as CO-GNN, NSD, and BuNN (Bamberger et al., 2025). Results for BuNN on Squirrel and Chameleon are not available, since we do not have access to their code.
For the remaining datasets, we compare against the classical GCN, GAT and SAGE; the models specifically tailored for heterophilic data GGCN (Yan et al., 2022), Geom-GCN (Pei et al., 2020), H2GCN (Zhu et al., 2020), GPRGNN (Chien et al., 2020), FAGCN (Bo et al., 2021) and MixHop (Abu-El-Haija et al., 2019); and the models GCNII (Chen et al., 2020) and PairNorm (Zhao and Akoglu, 2020) designed to alleviate oversmoothing. We use the 10 fixed splits proposed by Platonov et al. (2023) and Pei et al. (2020). We refer to Appendix˜B for further implementation details.
Results. Table˜1 and Table˜2 show that CSNN is the best-performing method in 9 out of 11 datasets. These results highlight our model’s capacity to deal with heterophilic graphs of different sizes and heterophily levels. We note CSNN often outperforms both NSD and CO-GNN by a significant margin.
6.3 Graph Classification
To assess the effectiveness of CSNNs on long-range tasks, we evaluate it on the peptides dataset from the Long Range Graph Benchmark Dwivedi et al. (2022). It is a dataset containing  15k graphs and two different tasks: peptides-func is a graph classification task, while peptides-struct is a regression one. We report average precision (AP) for peptides-func and mean absolute error (MAE) for peptides-struct.
Table 3: Performance comparison of models on the peptides datasets.
Model peptides-func $\uparrow$ peptides-struct $\downarrow$
GCN 68.60 ± 0.50 24.60 ± 0.07
GINE 66.21 ± 0.67 24.73 ± 0.17
GatedGCN 67.65 ± 0.47 24.77 ± 0.09
DReW 71.50 ± 0.44 25.36 ± 0.15
SAN 64.39 ± 0.75 25.45 ± 0.12
GPS 65.34 ± 0.91 25.09 ± 0.14
G-ViT 69.42 ± 0.75 24.49 ± 0.16
Exphormer 65.27 ± 0.43 24.81 ± 0.07
BuNN 72.76 ± 0.65 24.63 ± 0.12
CSNN 71.58 ± 0.80 24.32 ± 0.04
Setup. We follow the experimental setup of Tönshoff et al. (2024), and tune the network hyperparameters keeping the $\sim\!\!500k$ parameter budget proposed by Dwivedi et al. (2022) for fair comparison. We run the model using four different seeds and report mean and standard deviation of the evaluation metrics. The baselines are taken from Tönshoff et al. (2024) and we also compare against the results reported for BuNN by Bamberger et al. (2025).
Results. Our model achieves the best performance in the peptides-struct dataset, and second-best in the peptides-func, as shown in Table˜3. These results further strengthen our claims on the capacity of CSNNs to mitigate over-squashing and perform better in scenarios where long-range and under-reaching are known issues.
7 Conclusion
This work proposed Cooperative Sheaf Neural Networks, a novel SNN architecture that incorporates directionality in order to increase its efficiency by learning sheaves with conformal maps, allowing nodes to choose the optimal behavior in terms of information propagation with respect to its neighbors. We provided theoretical insights on how CSNN can alleviate over-squashing due to its capacity to smoothly modulate node behavior in information diffusion. We also validated its effectiveness on node and graph classification experiments on heterophilic graphs and long-range tasks.
Limitations and Future Works. While CSNN is not computationally more taxing than other SNNs, it is worth pointing that developing strategies to scale sheaf-based networks is a major research challenge. While we have used conformal maps to reduce the parameter complexity of restriction maps, we leave open the possibility that there are further ways to improve the scalability of CSNN. We also believe that efficient message-passing implementations could represent a step towards large-scale SNNs.
Another promising direction for future works is extending SNNs to cope with high-order structures like cell- and simplicial complexes, possibly allowing for more expressive models and promoting long-range communication with fewer layers.
8 Ethics and Reproducibility Statements
Ethics Statement. We do not foresee immediate negative societal or ethical impacts at this stage of the work.
Reproducibility Statement. Aiming to secure reproducibility of our work, we provide proofs of our theoretical results and in experiment detail in Appendix˜A and Appendix˜B. Moreover, we will provide a public code once the review process is complete.
Appendix A Proofs
A.1 Proof of Proposition 3.1
Proof.
Suppose $L_{\mathcal{F}}(\mathbf{X})_{i}$ does not depend of $\mathbf{x}_{j}$ for any $j$ neighbor of $i$. Since $L_{\mathcal{F}}(\mathbf{X})_{i}=\sum_{i,j\unlhd e}\mathcal{F}_{i\unlhd e}^{\top}\left(\mathcal{F}_{i\unlhd e}\mathbf{x}_{i}-\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}\right)$, this means $\mathcal{F}_{i\unlhd e}^{\top}\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}=0$. Therefore, $\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}\in\ker(\mathcal{F}_{i\unlhd e}^{\top})$, for any $j$ neighbor of $i$. Thus, $\mathcal{F}_{i\unlhd e}\mathbf{x}_{i}=0$ or $\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}=0$, for every $j$.
Note that, $\mathcal{F}_{i\unlhd e}\mathbf{x}_{i}=0$ implies that $L_{\mathcal{F}}(\mathbf{X})_{j}=\sum_{j,i\unlhd e}\mathcal{F}_{j\unlhd e}^{\top}\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}$.
If $\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}=0$ for every $j$, then $L_{\mathcal{F}}(\mathbf{X})_{j}=0$.
A.2 Proof of Proposition 4.1
@@ Appendix A Proofs
A.3 Proof of Proposition 4.2
@@
Let $t=1$, and fix a node $i$. Then we are essentially just using the composition described in Equation˜9 (up to normalization and learnable weights). In the equation we have a sum running over all neighbors $j$ of $i$ and another sum running over all neighbors $u$ of each $j$. So $u$ can be a $2$-hop neighbor of $i$ and we have that $i$ was updated with information from up to $2$-hop neighbors. Similarly, the node $u$ is updated by up to $2$-hop neighbors. Therefore, in the second layer $t=2$, $i$ was updated with information from up to $4$-hop neighbors.
@@
A.4 Proof of Proposition 4.3
@@
Choose a path from $i$ to $j$. So there are t1t-1 vertices between $i$ an $j$, say $v_{1},...,v_{t-1}$. In the first layer, let ${\mathbf{S}}_{j}$ and ${\mathbf{T}}_{v_{t-1}}$ be different of zero and all other source and target maps equal zero.
@@
Appendix B Experiment details